# Shifting the Lens: Detecting Malicious npm Packages using Large Language Models

### v2

@@ Abstract
Existing malicious code detection techniques can aid the manual review process by predicting which packages are likely to be malicious. However, these techniques often suffer from high misclassification rates. Therefore, malicious code detection techniques could be enhanced by adopting advanced, more automated approaches to achieve high accuracy and a low misclassification rate. The goal of this study is to assist security analysts in detecting malicious packages through the empirical study of using Large Language Models (LLMs) to detect malicious code in the npm ecosystem. We present SecurityAI, a malicious code review workflow to detect malicious code using ChatGPT. We leverage a benchmark dataset of 5,115 npm packages, of which 2,180 packages have malicious code.
We conducted a baseline comparison of GPT-3 and GPT-4 models with the state-of-the-art CodeQL static analysis tool, using 39 custom CodeQL rules developed in prior research to detect malicious Javascript code. We compare the effectiveness of static analysis as a pre-screener with SecurityAI workflow, measuring the number of files that need to be analyzed and the associated costs. Additionally, we performed a qualitative study to understand the types of malicious packages detected or missed by our workflow. Our baseline comparison demonstrates a 16% and 9% improvement over static analysis in precision and F1 scores, respectively. We attained precision and F1 scores of 91% and 94% for GPT-3, and 99% & 97% for GPT-4, respectively, with GPT-3 offering a cost-effective balance. Pre-screening files with a static analyzer reduces the number of files requiring LLM analysis by 77.9% and decreases costs by 60.9% for GPT-3 and 76.1% for GPT-4. Our qualitative analysis identified data theft, hidden backdoors, and suspicious domain connection categories as the top detected malicious packages. The lack of diversity in model-generated responses led to hallucinations, resulting in misclassification cases, with GPT-3 hallucinating more frequently.
@@ I Introduction
Open-source package repositories, like npm, PyPI, and RubyGems are often misused for software supply chain attacks [28]. Attackers attempt to execute supply chain attacks by injecting malicious code into existing packages or by creating new infected packages and tricking users into downloading them  [28, 36]. Given the increasing rate of software supply chain attacks, the demand for approaches to detect such attacks is also increasing. Unlike traditional security vulnerabilities that arise from accidental mistakes, malicious vulnerabilities involve attackers injecting intentional vulnerabilities into code. Classifying a package as malicious is challenging because malicious code detection necessitates comprehending the reasoning about the possible purpose and intent of the code.
The current malicious package detection [17, 19, 57, 35, 41, 42, 47, 46, 49] including static, rules, heuristics, differential analysis, and different machine learning techniques encounter challenges of incomplete solutions, high misclassification rates [37]. Therefore, these techniques aid in predicting the likelihood of a package being malicious or benign by providing additional information to the manual review process [37]. Additionally, conducting a comprehensive manual review to understand the code intent of a package is not scalable for software security analysists [8, 37]. An advanced detection approach that reduces the burden of manual review on security analysts and offers findings with minimal misclassification is needed. The goal of this study is to assist security analysts in detecting malicious packages through the empirical study of using Large Language Models (LLMs) to detect malicious code in the npm ecosystem.
Large Language Models (LLMs) are pre-trained models that can solve a variety of natural language processing (NLP) tasks by conditioning the model on appropriate prompts [10, 12, 40]. While the existing literature lacks the contribution of using LLMs for detecting malicious code, organizations are adopting LLMs as an advanced technique for threat analysis [4, 32]. However, assessing the performance of LLMs in production presents challenges: 1) LLMs overlooking actual malicious code; 2) validation due to the lack of ground truth; 3) after reporting potential malicious packages to the package registry, the package managers often only acknowledge receipt of the reports without confirming the presence of malicious code; and 4) comprehensive evaluation in a production environment is resource-intensive. Hence, we conduct a scientific study to evaluate LLMs’ performance in malicious code detection. Therefore, we address the following research questions (RQs)—
@@
  • RQ2: What is the effectiveness of using a static analyzer as a pre-screener with SecurityAI in terms of the number of files that need to be analyzed and the associated cost?
@@
In this study, we use ChatGPT models (GPT-3 and GPT-4) to evaluate the performance of LLMs in detecting malicious packages. LLMs based on the GPT (generative pre-trained transformer) architecture, particularly ChatGPT, have demonstrated superior performance in generating conclusions from new data using a model’s pre-trained knowledge [40, 60, 53, 31]. We leverage MalwareBench [56], a benchmark dataset, and studied 5,115 npm packages, including 2,180 malicious and 2,935 neutral packages.
We present SecurityAI, a malicious code review workflow to answer our research questions. The workflow leverages Iterative Self-Refinement [33] and Zero-Shot-Role-Play-Chain of Thought (CoT) prompting [25, 26] techniques in ChatGPT models to review Javascript packages. To the best of our knowledge, this research presents a first-ever study leveraging LLMs to detect malicious code using Iterative Self-Refinement and Zero-Shot-Role-Play-CoT prompting techniques and compare the effectiveness of the workflow with static analysis tools. To answer RQ1, The paper presents a baseline comparison between ChatGPT models and the mature, state-of-the-art commercial framework, CodeQL. We used 39 custom rules of CodeQL that have been used in prior study [16] to detect malicious Javascript packages. We compare results using the standard machine learning (ML) evaluation metrics, precision, recall, and F1 score. We answer RQ2 by comparing the effectiveness of using SecurityAI only versus using CodeQL as a pre-screener with SecurityAI based on the number of files analyzed and the associated costs. For RQ3, we provide qualitative insights by manually evaluating the model’s responses to understand the types of malicious packages identified or missed by ChatGPT and demonstrate the opportunities and strengths of these models. For the benefit of other researchers, we further report issues we encountered while using ChatGPT. Our contributions are:
  • SecurityAI, a malicious code review workflow to detect malicious packages in the npm ecosystem;
  • Malware-domain-specific prompt to trigger LLMs reasoning in detecting malicious code, available in Zenodo [3];
  • A baseline comparison of SecurityAI with CodeQL tool to understand the effectiveness of LLMs; and
@@
This paper is organized as follows: Section II includes our design approach, discussing our dataset, allocated cost, package selection, and model selection criteria. Section III covers our proposed workflow of SecurityAI. We answer our three RQs in section IV, section V, and section VI. Then, we discussed challenges in using LLM in section VII. The related work section is covered in section VIII. We close with limitations and discussion of our work (Section IX and X).
@@ II Dataset and LLM Selection
The section explores the dataset and model selection criteria. Section II-A discusses the benchmark dataset used in this study, Section II-B discusses LLMs selection process, and Section II-D explains our package selection criteria with cost constraint discussed in Section  II-C.
@@ II-A Benchmark Dataset: MalwareBench
Recent studies [36, 14, 20, 27] constructed malicious package datasets. However, these studies did not incorporate neutral package data into their labeled datasets. Neutral packages are needed to improve the accuracy of detection techniques. To address the lack of labeled neutral packages, Zahan et al. constructed MalwareBench [56], a labeled benchmark dataset of malicious and neutral npm packages. The benchmark dataset contained past malicious package dataset [36, 14, 20, 27] augmented by new industry data that includes malicious and neutral packages. Zahan et al. refer to a package as neutral when it has no discovered malicious code [56]. The ground-truth labels for the benchmark were determined using automatic and manual inspection. MalwareBench contains 3,523 malicious and 10,691 neutral npm packages. The malicious category comprises real-world malicious packages, while the neutral category includes popular, new, and randomly annotated real-world npm packages.
@@ II-C Cost Allocation
Due to the cost and time associated with using GPT models, we first allocated a budget for this project The budget will subsequently inform the scope of our analysis regarding the number of files or packages that can be assessed. We set up a budget of \$3000 because increasing the budget would extend the model’s computation time and the time needed for authors to manually review the model’s generated conclusions. We conducted a preliminary analysis of 1000 Javascript files using both GPT-3 and GPT-4 following the workflow mentioned in Section IV. We found that analyzing 1000 Javascript files with GPT-3 incurred a cost of \$10.5, while the same analysis with GPT-4 cost \$130. Armed with this cost information, we deduced that with a budget of \$3000, we could analyze around 20,000 files.
@@ III SecurityAI
The section discusses the multi-stage decision-maker workflow, SecurityAI , that leverages LLMs to detect malicious packages. OrgA (will disclosed upon paper acceptance) has been using SecurityAI for over a year to detect malicious packages within the npm, PyPI, Go, and Maven ecosystems. On average, SecurityAI is used to evaluate 17 million package versions quarterly in production. As of 2024, three authors are actively developing and maintaining the workflow. In Section III-A, III-B and III-C, we discuss our design decisions and the prompting techniques, and section III-D discuss our proposed SecurityAI workflow.
III-A Iterative Self-Refinement Approach
The SecurityAI workflow employs an iterative self-refinement [33] approach, systematically leveraging ChatGPT’s responses in a sequential analytical workflow. Iterative self-refinement is a process that involves creating an initial draft response generated by the model and subsequently refining the response based on the model’s self-provided feedback [33, 59]. Madaan et al. [33] showed that iterative self-refinement prompting had superior results compared to one-step response generation and improved performance (20%).
III-B Zero-Shot-Role-Play-CoT Prompting
We leveraged the Zero-Shot-CoT [25, 50, 54] technique along with Role-Play prompting [26] to enhance model reasoning in evaluating npm packages. Prior studies [25, 26, 54] showed that Zero-Shot-CoT and Role-Play prompting enhance model performance across multiple domains. Subsequently, in Role-Play prompting, the role provides context about the model’s identity and background, which serves as an implicit CoT trigger, thereby improving the quality of reasoning.
Figure 1: SecurityAI workflow
III-C Prompt Design
@@
We designed a two-fold prompt structure: the system role prompt serves as the instructions to set the context for how the model should behave. In contrast, the user role prompt comprises file content we want the models to review. Additionally, prior studies [6, 52, 51] demonstrate that augmenting the model with domain-specific information improves the model’s performance. Our prompting techniques were inspired by [61, 50, 54, 6, 25, 26, 54, 52, 51, 7], while also prioritizing a domain-targeted prompt using the author’s domain knowledge. We propose diverse augmentations of system role prompts to provide LLMs with more task-specific information. Table I, Table II and Table III contain the prompt used in this study.
Role Description. We defined different system roles for LLMs to follow in multiple steps (Section III-D) to improve accuracy. We assign the system role as SecureGPT, a JavaScript cybersecurity analyst and security researcher, to narrow the focus to JavaScript’s particular security issues. The system role contextualization provides background information and guides the model’s evaluation parameters, aligning the analysis with the defined scenario.
Task & Guidelines. We first augment the prompt by defining the task of analyzing JavaScript code for supply chain attacks, malicious behavior, and other security risks. Then, we provide guidelines to spot anomalies, such as hardcoded credentials, backdoors, and unusual behaviors, while noting code observations and avoiding unnecessary warnings.
Malware, Security Risk, and Obfuscated Code. We provide definitions and examples for malware, security risks, and obfuscated code. The targeted information helps the model recognize and differentiate between malware and other security vulnerabilities. Additionally, in malware detection, one of the primary challenges is prioritizing and differentiation between malware cases and security vulnerabilities. We address this challenge by incorporating scoring metrics — malware intent, security risk, and confidence — to standardize the evaluation output, enabling the model to quantify the severity of findings and assisting human reviewers in prioritizing potentially malicious files. Since attackers commonly use obfuscation to avoid detection, it is included to identify such cases. The confidence score also aids in evaluating the reliability of the model’s findings, while malware, security risk, and obfuscation are included to understand the nature of the code. Each scoring is defined in a range of 0-1, which we divide into different groups. Table I (D7-D9) are examples of scoring prompts.
Code Review & JSON Response. Motivated by [6, 52, 51], we included code-review prompts (Table I-D10) that guide the model in step-by-step thinking and generating conclusions by detecting key identifiers of code such as sources, sinks, flows, anomalies. Additionally, large-scale analysis in production requires machine-readable reports. Therefore, we improved the prompting technique by instructing the model to generate the output report via a JSON format (Table I-D11). Besides the report being machine-readable, following such templates forces the model to rethink its evaluation, conducting a sequential analysis (purpose-sources-sinks-flows-analysis-conclusion) and generating analysis reports to the specified format.
@@
User Role Prompt. The user role prompt comprises the content (Javascript files) we want the model to analyze. For each file, we include the code as a user prompt, following the prompt— Analyze the above code for malicious behavior. Remember to respond in the required JSON format. Consider ALL of the code carefully. Check the beginning, middle, and end of the code. Work step-by-step to get the right answer.
@@
III-D Proposed Workflow
Our SecurityAI workflow consists of four steps: the code as a user input and system role prompt; refining the responses in the next two steps; and evaluating with models generating a final analysis (Figure 1). Initially, we tuned the model to explore a wide range of ideas and possibilities. Later, the focus transitioned towards enhancing the model’s ability to refine its own responses, aiming to produce more precise analyses.
III-D1 Step 1: Initial Reports
@@
III-D2 Step 2: Critical Reports
@@
The critical reports step is designed to reassess the model’s initial report responses using system prompts and user inputs to perform an in-depth evaluation. As input, the system prompt (Table II) instructs the model, acting as a security researcher, to review and evaluate the initial findings thoroughly. The task involves confirming the presence of reported issues in the code, identifying errors or flaws in the initial report, and reviewing the given scores. The model is asked to justify any risk score above 0.5, ensure the score accurately reflects malicious behavior, challenge assumptions about vulnerabilities, and maintain consistency in scoring. For the user role prompt, we input the model-generated responses from Step 1 (section III-D1) and the file to be evaluated. The model generates a predetermined number of output reports (5 for GPT-3 and 3 for GPT-4). The model is tuned to a temperature parameter of 0.75 and a $top_{p}$ value of 0.6 to constrain the idea generation.
III-D3 Step 3: Final Report

### v3

# Leveraging Large Language Models to Detect npm Malicious Packages

@@ Abstract
Existing malicious code detection techniques demand the integration of multiple tools to detect different malware patterns, often suffering from high misclassification rates. Therefore, malicious code detection techniques could be enhanced by adopting advanced, more automated approaches to achieve high accuracy and a low misclassification rate. The goal of this study is to aid security analysts in detecting malicious packages by empirically studying the effectiveness of Large Language Models (LLMs) in detecting malicious code. We present SocketAI, a malicious code review workflow to detect malicious code. To evaluate the effectiveness of SocketAI, we leverage a benchmark dataset of 5,115 npm packages, of which 2,180 packages have malicious code. We conducted a baseline comparison of GPT-3 and GPT-4 models with the state-of-the-art CodeQL static analysis tool, using 39 custom CodeQL rules developed in prior research to detect malicious Javascript code. We also compare the effectiveness of static analysis as a pre-screener with SocketAI workflow, measuring the number of files that need to be analyzed and the associated costs. Additionally, we performed a qualitative study to understand the types of malicious activities detected or missed by our workflow. Our baseline comparison demonstrates a 16% and 9% improvement over static analysis in precision and F1 scores, respectively. GPT-4 achieves higher accuracy with 99% precision and 97% F1 scores, while GPT-3 offers a more cost-effective balance at 91% precision and 94% F1 scores. Pre-screening files with a static analyzer reduces the number of files requiring LLM analysis by 77.9% and decreases costs by 60.9% for GPT-3 and 76.1% for GPT-4. Our qualitative analysis identified data theft, suspicious domain connection, and arbitrary code execution as the top detected malicious activities.
@@ I Introduction
Open-source package repositories are often misused for software supply chain attacks [23]. Attackers attempt to execute supply chain attacks by injecting malicious code into existing packages or by creating new infected packages and tricking users into downloading them  [23, 32]. Classifying a package as malicious is challenging because malicious code detection necessitates comprehending the reasoning about the possible purpose and intent of the code. Given the increasing rate of software supply chain attacks, the demand for approaches to detect such attacks is also increasing.
The current malicious package detection [12, 14, 56, 31, 38, 39, 45, 44, 47] including static, rules, heuristics, differential analysis, and different machine learning techniques encounter challenges of incomplete solutions, high misclassification rates [33]. Additionally, no single technique can comprehensively address different attack types, like detecting malicious code injection, dependency confusion, and typosquat, and often fails to support languages like TypeScript and CoffeeScript or research packages [33]. As a result, practitioners need to integrate multiple tools to detect malware patterns. An advanced detection technique is needed to identify diverse attack types effectively and minimize manual review efforts with minimal misclassification. LLMs have the potential to address these limitations by detecting diverse attacks through their advanced understanding of context, semantics, and patterns, thereby reducing reliance on multiple tools.
While the existing literature lacks the contribution of using LLMs to detect malicious code, organizations are adopting LLMs as an advanced technique for threat analysis [3, 27]. However, assessing the performance of LLMs in production presents challenges: 1) comprehensive evaluation in a production environment is resource-intensive: 2) LLM missing actual malicious code (false negative); 2) validation due to the lack of ground truth; and 4) after reporting potential malicious packages to the package registry, the package managers often only acknowledge receipt of the reports without confirming the presence of malicious code. Hence, The goal of this study is to aid security analysts in detecting malicious packages by empirically studying the effectiveness of Large Language Models (LLMs) in detecting malicious code. In this study, we address the following research questions (RQs)—
@@
  • RQ2: What is the effectiveness of using a static analyzer as a pre-screener with SocketAI in terms of the number of files that need to be analyzed and the associated cost?
@@
To the best of our knowledge SocketAI, a malicious code review workflow is a first-ever study leveraging LLMs to detect malicious code and compares the effectiveness of the workflow with static analysis tools. The workflow leverages Iterative Self-Refinement [28], LLM-as-a-Judge [58] and Zero-Shot-Role-Play-Chain of Thought (CoT) prompting [20, 21] techniques in ChatGPT models (GPT-3 and GPT-4). We leveraged MalwareBench [55], a benchmark dataset, and studied 5,115 npm packages, including 2,180 malicious and 2,935 neutral packages, to evaluate the performance of LLMs.
To answer RQ1, The paper presents a baseline comparison between ChatGPT models and the mature, state-of-the-art commercial tool, CodeQL. We used 39 custom rules of CodeQL that have been used in prior study [11] to detect malicious Javascript packages. We compare results using the standard machine learning (ML) evaluation metrics, precision, recall, and F1 score. We answer RQ2 by comparing the effectiveness of using “SocketAI only” versus using “CodeQL as a pre-screener with SocketAI ” based on the number of files analyzed and the associated costs. For RQ3, we provide qualitative insights by manually evaluating the model’s responses to understand the types of malicious packages identified or missed by ChatGPT and demonstrate the opportunities and strengths of these models. For the benefit of other researchers, we further report issues we encountered while using ChatGPT. Our contributions are:
  • SocketAI, a malicious code review workflow to detect malicious packages in the npm ecosystem;
  • Malware-domain-specific prompt to trigger LLMs reasoning in detecting malicious code;
  • A baseline comparison of SocketAI with CodeQL tool to understand the effectiveness of LLMs; and
@@
This paper is organized as follows: Section II includes our design approach, discussing our dataset, allocated cost, package selection, and model selection criteria. Section III covers our proposed workflow of SocketAI. We answer our three RQs in section IV, section V, and section VI. Then, we discussed challenges in using LLM in section VII-B. The related work section is covered in section VIII. We close with limitations and discussion of our work (Section VII-C and IX).
@@ II Dataset and LLM Selection
Section II-A discusses the benchmark dataset used in this study, Section II-B discusses LLMs selection process, and Section II-D explains our package selection criteria with cost constraint discussed in Section  II-C.
@@ II-A Benchmark Dataset: MalwareBench
Recent studies [32, 9, 15, 22] constructed malicious package datasets. However, these studies did not incorporate neutral package data into their labeled datasets. Neutral packages are needed to improve the accuracy of detection techniques. To address the lack of labeled neutral packages, Zahan et al. constructed MalwareBench [55], a labeled benchmark dataset of malicious and neutral npm packages. Zahan et al. refer to a package as neutral when it has no discovered malicious code [55]. The ground-truth labels for the benchmark were established using a hybrid approach: researchers manually reviewed and confirmed all malicious cases flagged by automated tools, while neutral cases were initially identified through automation and further validated via manual analysis of randomly selected samples. MalwareBench contains 3,523 malicious and 10,691 neutral npm packages. The malicious category comprises real-world malicious packages, while the neutral category includes popular, new, and randomly annotated real-world npm packages.
@@ II-C Cost Allocation
Due to the cost and time associated with using GPT models, we first allocated a budget for this project. The budget will inform the scope of our analysis regarding the number of files or packages that can be assessed. We set up a budget of \$3000 because increasing the budget would extend the model’s computation time and the time needed for authors to manually review the model’s generated response. We conducted a preliminary analysis of 1000 Javascript files using both GPT-3 and GPT-4 following the workflow mentioned in Section IV. We found that analyzing 1000 Javascript files with GPT-3 incurred a cost of \$10.5, while the same analysis with GPT-4 cost \$130. Armed with this cost information, we deduced that with a budget of \$3000, we could analyze around 20,000 files.
@@ III SocketAI
The section discusses the multi-stage decision-maker workflow of SocketAI. Socket has been using SocketAI to detect malicious packages in production for npm, PyPI, Go, and Maven ecosystems. On average, SocketAI is used to evaluate 17 million package versions quarterly in production. As of 2024, three authors are actively developing and maintaining the workflow. In Section III-A, and III-B, we discuss our design decisions and the prompting techniques, and section III-C discusses our proposed SocketAI workflow.
III-A Leveraging LLM-Based Techniques
@@
Since SocketAI is used in production to review different ecosystems, we designed our domain-specific prompt to be model-agnostic and adaptable to open and closed-source models where the prompt techniques are fundamental to any LLM. Our prompting techniques were inspired by advanced LLM-based techniques by [28, 58, 20, 48, 52], while also prioritizing a domain-targeted prompt using the author’s domain knowledge. Prior studies [5, 50, 49] demonstrate that augmenting the model with domain-specific information improves the model’s performance.
The SocketAI workflow employs an iterative self-refinement [28] approach, systematically leveraging LLM’s responses in a sequential analytical workflow. Iterative self-refinement is a process that involves creating an initial draft response generated by the model and subsequently refining the response based on the model’s self-provided feedback [28, 57]. Madaan et al. [28] showed that iterative self-refinement prompting had superior results compared to one-step response generation and improved performance (20%). Then, we also leveraged the LLM-as-a-Judge approach [58] to evaluate and score the outputs against specific criteria. Self-refinement is a loop of generation and improvement, while LLM-as-a-Judge is better for scoring or verifying outputs against criteria. We used the Zero-Shot-CoT [20, 48, 52] technique along with Role-Play prompting [21] to enhance model reasoning in evaluating npm packages. Prior studies [20, 21, 52] showed that Zero-Shot-CoT and Role-Play prompting enhance model performance across multiple domains. Subsequently, in Role-Play prompting, the role provides context about the model’s identity and background, which serves as an implicit CoT trigger, thereby improving the quality of reasoning.
Figure 1: SocketAI workflow
III-B Prompt Design
@@
We designed a two-fold prompt structure: the system role prompt serves as the instructions to set the context for how the model should behave. We assign different system roles for LLMs to follow in multiple steps (Section III-C) such as SecureGPT, a JavaScript cybersecurity analyst and security researcher, to narrow the focus to JavaScript’s particular security issues. The user role prompt comprises the input content (npm packages files) we want the models to review. For each file, we include the code as a user prompt, following the prompt— Analyze the above code for malicious behavior. Remember to respond in the required JSON format. Consider ALL of the code carefully. Check the beginning, middle, and end of the code. Work step-by-step to get the right answer. Table I, Table II and Table III contain the prompts used in this study.
Task & Guidelines. We first augment the prompt by defining the task of reviewing JavaScript code for supply chain attacks, malicious behavior, and other security risks. Then, we provide guidelines to spot anomalies, such as hardcoded credentials, backdoors, and unusual behaviors, while noting code observations and avoiding unnecessary warnings.
Malware, Security Risk, and Obfuscated Code. We provide definitions and examples for malware, security risks, and obfuscated code. The targeted information helps the model recognize and differentiate between malware and other security vulnerabilities. Additionally, in malware detection, one of the primary challenges is prioritizing and differentiation between malware cases and security vulnerabilities. Motivated by LLM-as-a-Judge approach [58], we address this challenge by incorporating scoring metrics— malware intent, security risk, and confidence — to standardize the evaluation output, enabling the model to quantify the severity of findings and assisting human reviewers in prioritizing potentially malicious files. Since attackers commonly use obfuscation to avoid detection, it is included to identify such cases. The confidence score also aids in evaluating the reliability of the model’s findings, while malware, security risk, and obfuscation are included to understand the nature of the code. Each scoring is defined in a range of 0-1, which we divide into different groups. Table I (D7-D9) are examples of scoring prompts.
Requirement: Code Review & JSON Response. Motivated by [5, 50, 49], we included code-review prompts (Table I-D10) that guide the model in step-by-step thinking and generating conclusions by detecting key identifiers of code such as sources, sinks, flows, anomalies. Additionally, large-scale analysis in production requires machine-readable reports. Therefore, we improved the prompting technique by instructing the model to conduct a sequential analysis (purpose-sources-sinks-flows-analysis-conclusion) and to generate the output report via a JSON format (Table I-D11). Besides the report being machine-readable, following such templates forces the model to rethink its evaluation and generate analysis reports in the specified format.
@@
III-C Proposed Workflow
Our SocketAI workflow comprises four steps (Figure 1): code as user input along with a system role prompt (Step 1); refining the model’s responses through iterative self-refinemnet in the next two steps (Steps 2 and 3); and finally, adding humans in the loop to evaluate the model-generated report (Step 4).
III-C1 Step 1: Initial Reports
@@
III-C2 Step 2: Critical Reports
@@
The critical reports step is designed to reassess the model’s initial report responses to perform an in-depth evaluation. As input, the system prompt (Table II) instructs the model, acting as a security researcher, to review and evaluate the initial findings thoroughly. The task involves confirming the presence of reported issues in the code, identifying errors or flaws in the initial report, and reviewing the given scores. The model is asked to justify any risk score above 0.5, ensure the score accurately reflects malicious behavior, challenge assumptions about vulnerabilities, and maintain consistency in scoring. For the user role prompt, we input the model-generated responses from Step 1 (section III-C1) and the file to be evaluated. The model generates a predetermined number of output reports (5 for GPT-3 and 3 for GPT-4). The model is tuned to a temperature parameter of 0.75 and a $top_{p}$ value of 0.6 to constrain the idea generation.
III-C3 Step 3: Final Report